# Causal Classification of Pathological Misner-Type Spacetimes

### v2

@@ 6.3. Pseudo-Reissner-Nordström timelike geodesics
[…]
@@
Due to the geodesic incompleteness, the embedded pseudo-Reissner-Nordström cylinder with $\theta=\textrm{const}$ and $\phi=\textrm{const}$, restricted to $M_{1}\cup M_{2}$, resembles the situation in the $2$-dimensional Misner space and cylindrical pseudo-Schwarzschild spacetime.
@@ 7. Conclusion
We have investigated a family of causality-violating spacetimes—specifically, the pseudo-Schwarzschild, pseudo-Reissner-Nordström, and Misner spacetimes—that share fundamental causal structures, including the presence of chronology horizons, Cauchy horizons, radial geodesics, and acausal regions. All these spacetimes can be expressed in a warped product form, with metrics decomposing into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$. This structure allows us to focus on their $2$-dimensional cylindrical analogues. In Eddington–Finkelstein coordinates, these cylindrical versions can be brought to a canonical form $ds^{2}=g_{\nu\nu}d\nu^{2}+2g_{\nu r}d\nu dr$, from which the radial geodesic equations (Section 3) can be systematically derived. Renaming the coordinates as $\nu=\varphi$ and $r=T$, the connection to Misner spacetime becomes manifest.
@@
In each case, the base manifold $M_{Z}$ of the warped product structure is conformally related to a region of Minkowski spacetime and admits a Killing vector field whose orbits closely resemble those of the boost-generating vector field $X=x\,\partial_{t}+t\,\partial_{x}$ in two-dimensional Minkowski space. This Killing vector field generates a one-parameter group $G$ of isometries, some of whose orbits are timelike. A discrete subgroup $G_{0}\subset G$ acts properly and freely on $M_{Z}$, resulting in a Misner-like structure on the quotient $M_{Z}/G_{0}$, where the timelike orbits of $G$ project to closed timelike curves. [^11] As described in the classic book by Hawking and Ellis [11], Misner space is constructed by taking a quotient of half of Minkowski spacetime under a discrete subgroup of the one-parameter isometry group generated by the vector field $X=x\,\partial_{t}+t\,\partial_{x}$, which acts properly and freely in that region. A similar isometry group orbit structure appears in the $(t,r)$-planes of both the Schwarzschild and Reissner-Nordström spacetimes, allowing for analogous quotient constructions in these cases as well. From a mathematical standpoint, this conformal structure renders the similarities in causal properties relatively unsurprising, as they directly reflect known features of Misner spacetime. Nevertheless, from a physics perspective, the results are noteworthy. The spacetimes examined differ markedly in their physical interpretation and geometric origin—ranging from a flat vacuum solution of the Einstein equations, to an asymptotically flat black hole solution of the vacuum Einstein equations, to a model that violates the weak energy condition and requires ‘exotic matter’ as its source—yet they exhibit analogous causal behavior. Understanding these similarities across physically distinct settings provides valuable insights into how causality can be broken or preserved under varying geometrical and physical conditions.
@@
In the context of ongoing research on spacetimes with closed timelike curves [1, 4, 9, 19], our findings offer a unified framework to view several causality-violating spacetimes through a common geometric and causal lens. The work reveals the intricate relationships between these examples and motivates a broader investigation into the interplay between geometry, causality, and spacetime topology. For instance, it would be compelling to investigate whether other spacetime models, such as the pseudo-Kerr spacetime [3]—which is a rotating generalization of the pseudo-Schwarzschild metric and describes the spacetime around a rotating, uncharged, axially symmetric black hole —also belong to this family of causality-violating, Misner-type spacetimes.
Finally, this article proposes a conjecture concerning the isocausality of the spacetimes under consideration. A rigorous proof of the conjecture, as well as a deeper investigation into its physical implications, is left for future work. We view this article as a stepping stone toward a broader classification of causality-violating spacetimes. The ansatz presented here could be extended to more complex models, where tools such as isocausality may offer a systematic framework for organizing, comparing, and understanding their causal structures.

### v3

@@ 6.3. Pseudo-Reissner-Nordström timelike geodesics
(6.4) $\nu(r)=\int\frac{\frac{-1}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}\left(\xi-\omega\sqrt{\xi^{2}-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}\right)}{\omega\sqrt{\xi^{2}-(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}}dr.$
@@
$\nu(r)=\int\frac{-1}{(\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1)}\left(\omega-1\right)dr.$
Due to the geodesic incompleteness, the embedded pseudo-Reissner-Nordström cylinder with $\theta=\textrm{const}$ and $\phi=\textrm{const}$, restricted to $M_{1}\cup M_{2}$, resembles the situation in the 22–dimensional Misner space and cylindrical pseudo-Schwarzschild spacetime.
7. Isocausality
In this section we investigate the causal equivalence properties—in the sense of isocausality as introduced by García-Parrado and Senovilla—between the three model spacetimes considered in this work: the Misner spacetime $(\mathcal{M}_{\mathrm{M}},g_{\mathrm{M}})$, the pseudo-Schwarzschild spacetime $(\mathcal{M}_{\mathrm{PS}},g_{\mathrm{PS}})$, and the pseudo-Reissner-Nordström spacetime $(\mathcal{M}_{\mathrm{PRN}},g_{\mathrm{PRN}})$.
Definition 1 (Isocausality).
Two time-oriented spacetimes $(\mathcal{M},g)$ and $(\mathcal{N},h)$ are isocausal if there exist smooth maps
$\Phi:\mathcal{M}\longrightarrow\mathcal{N},\qquad\Psi:\mathcal{N}\longrightarrow\mathcal{M}$
such that both $\Phi$ and $\Psi$ are causal, i.e. their differentials map future-directed causal vectors to future-directed causal vectors. No bijectivity or inverse relationship between $\Phi$ and $\Psi$ is required by this definition.
In several results below we prove a stronger property: the causal maps $\Phi$ and $\Psi$ are actually mutual diffeomorphisms (smooth bijections with smooth inverses that are also causal) on the universal covers and on certain causally regular regions of the spacetimes. For the compactified models, global mutual diffeomorphisms exist only under an explicit equivariance condition relating the deck group actions; without this condition, isocausality in the strict sense is generally not guaranteed.
Each of our model spacetimes is a warped product
$\mathcal{M}_{i}\;=\;\mathcal{M}^{(Z)}_{i}\times_{r}H^{2}$
with 22–dimensional Lorentzian base $(\mathcal{M}^{(Z)}_{i},g^{(Z)}_{i})$ admitting coordinates $(\nu,r)$ in which the metric takes the canonical Eddington-Finkelstein-type form
(7.1) $g^{(Z)}_{i}\;=\;f_{i}(r)\,d\nu^{2}+2\,d\nu\,dr,$
and $H^{2}$ the hyperbolic space of constant negative curvature endowed with its standard metric $g_{H^{2}}$. The explicit functions $f_{i}(r)$ for $i=\mathrm{M},\mathrm{PS},\mathrm{PRN}$ have been given in the preceding sections.
Because all three bases share the same canonical structure (7.1), one expects the existence of explicit coordinate transformations mapping each $g^{(Z)}_{i}$ to any other $g^{(Z)}_{j}$ up to a smooth, strictly positive conformal factor. Such transformations preserve the null cones on the base and, when extended trivially over $H^{2}$, preserve the full causal structure of the warped product. This observation motivates the main results of this section:
  • In Proposition 7.1 we construct explicit causal bijections between the universal covers of the three models, proving that they are pairwise isocausal in the covering space. We also identify causally regular regions in the original spacetimes on which the same holds.
  • Part (C) of Proposition 7.1 provides a clean necessary and sufficient condition (equivariance with respect to the deck group actions and single-valuedness of the conformal factor) under which the isocausality descends to the compactified spacetimes.
  • Corollary 1 clarifies that, in the absence of such equivariance, one obtains only a one-way causal relation between the compactified models.
  • Remark 7.2 emphasizes that this one-way causal relation should be regarded as the generic global situation unless the deck-group matching is verified explicitly.
We now turn to the formal statements and proofs.
Proposition 7.1 (Isocausality of the Misner-type family).
Let $(\mathcal{M}_{\mathrm{M}},g_{\mathrm{M}})$ be Misner spacetime, $(\mathcal{M}_{\mathrm{PS}},g_{\mathrm{PS}})$ the pseudo-Schwarzschild spacetime, and $(\mathcal{M}_{\mathrm{PRN}},g_{\mathrm{PRN}})$ the pseudo-Reissner-Nordström spacetime, with the $1+1$ radial (base) metrics written in the canonical Eddington-Finkelstein-type form
(7.2) $g^{(Z)}\;=\;f(r)\,d\nu^{2}+2\,d\nu\,dr,$
where, on the base coordinates used in this paper,
$f_{\mathrm{M}}(T)=T,\qquad f_{\mathrm{PS}}(r)=-\Big{(}\frac{2m}{r}-1\Big{)},\qquad f_{\mathrm{PRN}}(r)=-\Big{(}\frac{2m}{r}-\frac{q^{2}}{r^{2}}-1\Big{)},$
and the full 44–metrics are the warped products $g_{i}=g^{(Z)}_{i}\oplus r^{2}g_{H^{2}}$ with the hyperbolic fiber $H^{2}$. Denote universal covers by $(\widetilde{\mathcal{M}}_{i},\tilde{g}_{i})$ and the covering projections by $\pi_{i}$.
(A) Local / covering isocausality. For every pair $(i,j)\in\{\mathrm{M},\mathrm{PS},\mathrm{PRN}\}^{2}$ there exists a smooth bijection
$\widetilde{\Phi}_{ij}:\widetilde{\mathcal{M}}_{i}\longrightarrow\widetilde{\mathcal{M}}_{j}$
whose differential preserves future-directed causal vectors and whose inverse $\widetilde{\Phi}_{ji}=\widetilde{\Phi}_{ij}^{-1}$ has the same property. In particular the universal covers $(\widetilde{\mathcal{M}}_{i},\tilde{g}_{i})$ are pairwise isocausal in the García-Parrado-Senovilla sense.
(B) Isocausality on causally regular regions. Let $U_{i}\subset\mathcal{M}_{i}$ denote open regions obtained by removing (a) neighborhoods of curvature singularities, and (b) neighborhoods of null hypersurfaces where $f_{i}=0$, so that $f_{i}$ is smooth and nowhere zero on $U_{i}$ and the $S^{1}$–orbits (if present) can be unwrapped. Then for every pair $(i,j)$ there exists a diffeomorphism $\Phi_{ij}:U_{i}\to U_{j}$ such that both $\Phi_{ij}$ and $\Phi_{ij}^{-1}$ map future-directed causal vectors to future-directed causal vectors; i.e. $(U_{i},g_{i})$ and $(U_{j},g_{j})$ are isocausal.
(C) Global isocausality (quotient) criterion. If the covering-space diffeomorphism $\widetilde{\Phi}_{ij}$ constructed in (A) is equivariant with respect to the deck groups (periodic identifications) $\Gamma_{i},\Gamma_{j}$—meaning there exists $k\in\{\pm 1\}$ such that
$\widetilde{\Phi}_{ij}\circ\gamma=\gamma^{k}\circ\widetilde{\Phi}_{ij}\quad\text{for all }\gamma\in\Gamma_{i},$
and if the induced conformal factor descends to a smooth, strictly positive single-valued function on the quotient—then $\widetilde{\Phi}_{ij}$ descends to a smooth map
$\Phi_{ij}:\mathcal{M}_{i}\longrightarrow\mathcal{M}_{j}$
that is a global causal bijection with causal inverse. In this case $(\mathcal{M}_{i},g_{i})$ and $(\mathcal{M}_{j},g_{j})$ are globally isocausal. If the equivariance condition fails (periods mismatch, or conformal factor not single-valued), the covering-space result does not imply global isocausality; at most a one-way causal relation $\mathcal{M}_{i}\prec\mathcal{M}_{j}$ can be guaranteed.
Proof.
We prove (A)–(C) by an explicit local construction on the $1+1$ bases and then extend trivially on the hyperbolic fiber.
Step 1: choice of domains. Fix any pair $(i,j)$ and choose open domains on the corresponding bases where $f_{i}$ and $f_{j}$ are smooth and nonvanishing (exclude neighborhoods of roots of $f$ and curvature singularities). On these domains the metrics are nondegenerate and causal cones are well defined.
Step 2: tortoise coordinate and model map. For each model define a (local) tortoise coordinate $r^{\ast}_{i}$ by any smooth antiderivative of
$\frac{dr}{dr^{\ast}_{i}}\;=\;\frac{|f_{i}(r)|}{2},$
so that $dr^{\ast}_{i}=(2/|f_{i}|)\,dr$. (Any smooth choice suffices on the fixed domain.) Fix a nonzero constant $\alpha\in\mathbb{R}$ and define a map on the $1+1$ base by
(7.3) $(\nu,r)\longmapsto(\phi,R)=\Big{(}\phi=\alpha\,\nu,\;R=\exp\!\big{(}\alpha\,r^{\ast}_{i}(r)\big{)}\Big{)}.$
Extend this map trivially on the hyperbolic fiber $(\theta,\varphi)\mapsto(\theta,\varphi)$ to obtain a map $\widetilde{\Phi}_{ij}$ between the corresponding covered regions of $\widetilde{\mathcal{M}}_{i}$ and $\widetilde{\mathcal{M}}_{j}$.
Step 3: pullback computation and conformal factor. A direct computation (cf. the calculation in Sec. 5.1.1) gives, on the base,
$\widetilde{\Phi}_{ij}^{*}\!\bigl{(}f_{j}(R)\,d\phi^{2}+2\,d\phi\,dR\bigr{)}=\Omega_{ij}(r)\,\bigl{(}f_{i}(r)\,d\nu^{2}+2\,d\nu\,dr\bigr{)},$
where the factor $\Omega_{ij}(r)$ is a smooth function determined by $f_{i},f_{j},\alpha$ and $r^{\ast}_{i}(r)$. Because $R>0$ and $r^{\ast}_{i}$ is smooth, one checks that $\Omega_{ij}(r)>0$ on the chosen domain (in particular one may choose $\alpha$ so that $\Omega_{ij}$ is positive everywhere on the domain). Thus the pullback equals the source metric times a strictly positive conformal factor.
Conformal rescalings preserve null cones in $1+1$ dimensions (and more generally they preserve the causal character of vectors), hence $d\widetilde{\Phi}_{ij}$ maps future-directed causal vectors of $\tilde{g}_{i}$ to future-directed causal vectors of $\tilde{g}_{j}$. The inverse map is obtained by reversing the change of variables: given $(\phi,R)$, recover $r^{\ast}_{j}=\alpha^{-1}\ln R$, then invert $r^{\ast}_{j}\mapsto r$ and set $\nu=\alpha^{-1}\phi$. The same computation applied in the reverse direction produces a smooth positive conformal factor for $\widetilde{\Phi}_{ji}$, so the inverse is also causal. This proves (A): the universal covers admit mutual causal bijections and are therefore pairwise isocausal.
Step 4: extension to warped product and causally regular regions. The full 44–metric is the warped product $g_{i}=g^{(Z)}_{i}\oplus r^{2}g_{H^{2}}$. Since the fiber metric $g_{H^{2}}$ is Riemannian (positive definite), the causal sign of a tangent vector $X=(X_{Z},X_{H})$ is determined by the base component up to the positive contribution of the fiber weighted by $r^{2}$. With $\widetilde{\Phi}_{ij}^{*}(g^{(Z)}_{j})=\Omega_{ij}\,g^{(Z)}_{i}$ and the fiber left unchanged (or mapped identically), one has
$\widetilde{\Phi}_{ij}^{*}(g_{j})(X,X)=\Omega_{ij}\,g^{(Z)}_{i}(X_{Z},X_{Z})+\bigl{(}r_{j}\circ\widetilde{\Phi}_{ij}\bigr{)}^{2}g_{H^{2}}(X_{H},X_{H}).$
Because $\Omega_{ij}>0$ and $g_{H^{2}}$ is positive definite, the causal cone on the pulled-back metric coincides with that of $g_{i}$ up to the positive scaling; therefore $\widetilde{\Phi}_{ij}$ preserves causal directions on the full 44–space. Restricting to quotient regions $U_{i}\subset\mathcal{M}_{i}$ where periodic identifications have been removed (or after passing to universal covers) gives the statement (B).
Step 5: descent / equivariance criterion (C). The maps above are constructed on universal covers. The covering projections $\pi_{i}:\widetilde{\mathcal{M}}_{i}\to\mathcal{M}_{i}$ have deck groups $\Gamma_{i}\cong\mathbb{Z}$ generated by the periodic identification in the angular/cylindrical coordinate. If $\widetilde{\Phi}_{ij}$ satisfies the equivariance condition $\widetilde{\Phi}_{ij}\circ\gamma=\gamma^{k}\circ\widetilde{\Phi}_{ij}$ for all $\gamma\in\Gamma_{i}$ and some fixed $k=\pm 1$, then $\widetilde{\Phi}_{ij}$ descends to a well defined smooth map $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$ with $\pi_{j}\circ\widetilde{\Phi}_{ij}=\Phi_{ij}\circ\pi_{i}$. Since $d\widetilde{\Phi}_{ij}$ preserves causal cones and the conformal factor is single-valued and strictly positive on the quotient by equivariance, the descended map $d\Phi_{ij}$ also preserves future-directed causal vectors. The inverse descends as well (equivariance for the inverse follows from equivariance of the original map), so $\Phi_{ij}$ is a global causal bijection with causal inverse: the compactified quotients are globally isocausal. If equivariance fails (periods mismatch or the conformal factor is multi-valued on the quotient) then the composition $\pi_{j}\circ\widetilde{\Phi}_{ij}\circ\pi_{i}^{-1}$ fails to be a bijection (or fails to be single-valued) and one obtains at best a one-way causal map $\mathcal{M}_{i}\to\mathcal{M}_{j}$ (hence $\mathcal{M}_{i}\prec\mathcal{M}_{j}$ but not necessarily the converse).
This completes the proof of (A)–(C). ∎
Corollary 1 (Default global outcome without equivariance).
Let $(\mathcal{M}_{i},g_{i})$ and $(\mathcal{M}_{j},g_{j})$ be any two of the Misner, pseudo-Schwarzschild, or pseudo-Reissner-Nordström spacetimes. Suppose there exists an isocausality diffeomorphism
$\widetilde{\Phi}_{ij}:\widetilde{\mathcal{M}}_{i}\longrightarrow\widetilde{\mathcal{M}}_{j}$
between their universal covers, as in Proposition 7.1(A), but that the equivariance hypothesis in Proposition 7.1(C) fails (e.g. the deck-group periods do not match up to an integer winding, or the conformal factor is not single-valued on the quotient). Then the descended map
$\Phi_{ij}\;:=\;\pi_{j}\circ\widetilde{\Phi}_{ij}\circ\pi_{i}^{-1}\;:\;\mathcal{M}_{i}\longrightarrow\mathcal{M}_{j}$
is a well-defined smooth causal map (future-directed causal vectors are mapped to future-directed causal vectors), but $\Phi_{ij}$ need not be bijective nor admit a causal inverse on the quotient. Thus, in the absence of deck-equivariance, one has only the one-way causal relation
$\mathcal{M}_{i}\prec\mathcal{M}_{j}$
globally, despite mutual isocausality at the covering level.
Proof.
Let $\pi_{i}:\widetilde{\mathcal{M}}_{i}\to\mathcal{M}_{i}$ and $\pi_{j}:\widetilde{\mathcal{M}}_{j}\to\mathcal{M}_{j}$ be the universal covering projections, and let $\widetilde{\Phi}_{ij}:\widetilde{\mathcal{M}}_{i}\to\widetilde{\mathcal{M}}_{j}$ be the causal diffeomorphism between covers whose existence is asserted in Proposition 7.1(A). Denote by $\Gamma_{i}$ and $\Gamma_{j}$ the corresponding deck groups. Recall that a map $\widetilde{\Phi}_{ij}$ descends to a well-defined map $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$ if and only if it is equivariant with respect to the deck actions, i.e. there exists $k\in\{\pm 1\}$ such that
$\widetilde{\Phi}_{ij}\circ\gamma=\gamma^{k}\circ\widetilde{\Phi}_{ij}\qquad\text{for all }\gamma\in\Gamma_{i}.$
If equivariance fails then there exists $\gamma\in\Gamma_{i}$ for which $\widetilde{\Phi}_{ij}\circ\gamma\neq\gamma^{k}\circ\widetilde{\Phi}_{ij}$ for every $k\in\{\pm 1\}$. Pick any $x\in\mathcal{M}_{i}$ and choose two distinct lifts $\tilde{x},\tilde{x}^{\prime}=\gamma\tilde{x}\in\widetilde{\mathcal{M}}_{i}$. Then
@@ 7. Isocausality
are the two candidate values for a putative descended map $\Phi_{ij}(x)$ obtained by composing the projections with $\widetilde{\Phi}_{ij}$. But, because equivariance fails, these two points need not coincide in $\mathcal{M}_{j}$, so the composition
$\pi_{j}\circ\widetilde{\Phi}_{ij}\circ\pi_{i}^{-1}$
is not single-valued on $\mathcal{M}_{i}$ and hence does not define a well-defined map $\Phi_{ij}:\mathcal{M}_{i}\to\mathcal{M}_{j}$.
Consequently, the cover construction alone does not produce a canonical causal map between the compactified spacetimes unless equivariance holds. Since isocausality of $(\mathcal{M}_{i},g_{i})$ and $(\mathcal{M}_{j},g_{j})$ requires the existence of causal maps in both directions on the quotients themselves, the failure of descent means that global isocausality cannot be concluded from the covering diffeomorphism without an independent argument that either (i) a descended causal map exists by some other construction, or (ii) the equivariance obstruction can be removed. This completes the proof. ∎
Remark 7.2 (On the generic global situation).
In the absence of an explicit verification of the deck-equivariance condition in Proposition 7.1(C), the generic expectation for the compactified Misner-type spacetimes considered here is that mutual isocausality will not hold globally, even though their universal covers are mutually isocausal. Without equivariance, any global causal relationship must be established by direct construction, and may exist in only one direction.
Physical meaning of global vs. covering isocausality. The distinction between isocausality at the level of the universal covers and the compactified spacetimes has a clear operational interpretation. On the universal cover, the causal structure is the “local” one perceived by any observer restricted to a simply connected region: the shape of the light cones, the accessibility of events along causal curves, and the arrangement of horizons are indistinguishable across the three models. In this sense, freely falling observers with access only to local measurements would not be able to tell which spacetime they inhabit; the causal physics is the same.
However, when passing to the compactified spacetimes by quotienting along closed orbits, additional global identifications introduce genuinely new causal phenomena, most notably the appearance and arrangement of closed timelike curves (CTCs). Failure of the deck-equivariance condition means that these identifications differ between the models: a causal curve that is closed in one model may fail to be closed in another, or the chronology-violating set may have a different extent. Operationally, an observer traversing a large loop along a timelike orbit in one spacetime might return to their starting event, while in another spacetime with the same local light cone structure they would not —or would reappear at a different event. Thus, the one-way causal relation $\mathcal{M}_{i}\prec\mathcal{M}_{j}$ reflects the fact that every causal path available in $\mathcal{M}_{i}$ is realisable in $\mathcal{M}_{j}$, but not necessarily the other way around. This is a genuinely global effect that cannot be detected from purely local measurements.
@@ 8. Conclusion
We have investigated a family of causality-violating spacetimes—specifically, the pseudo-Schwarzschild, pseudo-Reissner-Nordström, and Misner spacetimes—that share fundamental causal structures, including the presence of chronology horizons, Cauchy horizons, radial geodesics, and acausal regions. All these spacetimes can be expressed in a warped product form, with metrics decomposing into the cylindrical metric $g_{Z}$ and the hyperbolic metric $g_{H^{2}}$. This structure allows us to focus on their 22–dimensional cylindrical analogues. In Eddington-Finkelstein coordinates, these cylindrical versions can be brought to a canonical form $ds^{2}=g_{\nu\nu}d\nu^{2}+2g_{\nu r}d\nu dr$, from which the radial geodesic equations (Section 3) can be systematically derived. Renaming the coordinates as $\nu=\varphi$ and $r=T$, the connection to Misner spacetime becomes manifest.
@@
In each case, the base manifold $M_{Z}$ of the warped product structure is conformally related to a region of Minkowski spacetime and admits a Killing vector field whose orbits closely resemble those of the boost-generating vector field $X=x\,\partial_{t}+t\,\partial_{x}$ in two-dimensional Minkowski space. This Killing vector field generates a one-parameter group $G$ of isometries, some of whose orbits are timelike. A discrete subgroup $G_{0}\subset G$ acts properly and freely on $M_{Z}$, resulting in a Misner-like structure on the quotient $M_{Z}/G_{0}$, where the timelike orbits of $G$ project to closed timelike curves. [^11] As described in the classic book by Hawking and Ellis [11], Misner space is constructed by taking a quotient of half of Minkowski spacetime under a discrete subgroup of the one-parameter isometry group generated by the vector field $X=x\,\partial_{t}+t\,\partial_{x}$, which acts properly and freely in that region. A similar isometry group orbit structure appears in the (t,r)(t,r)–planes of both the Schwarzschild and Reissner–Nordström spacetimes, allowing for analogous quotient constructions in these cases as well. From a mathematical standpoint, this conformal structure renders the similarities in causal properties relatively unsurprising, as they directly reflect known features of Misner spacetime. Nevertheless, from a physics perspective, the results are noteworthy. The spacetimes examined differ markedly in their physical interpretation and geometric origin—ranging from a flat vacuum solution of the Einstein equations, to an asymptotically flat black hole solution of the vacuum Einstein equations, to a model that violates the weak energy condition and requires ‘exotic matter’ as its source—yet they exhibit analogous causal behavior. Understanding these similarities across physically distinct settings provides valuable insights into how causality can be broken or preserved under varying geometrical and physical conditions.
A central result of this work is the formal proof of the isocausality of the Misner-type family at the covering level. In Proposition 7.1 we showed that the universal covers of the Misner, pseudo-Schwarzschild, and pseudo-Reissner-Nordström spacetimes are pairwise isocausal via explicit smooth bijections whose differentials preserve and reflect all future-directed causal cones. This isocausality also holds on suitable open, causally regular regions of the compactified spacetimes where the chronology-violating sets and Cauchy horizons are excised. Furthermore, we identified an explicit and checkable deck-equivariance criterion under which the isocausality descends to the full compactified spacetimes. When this criterion fails, Corollary 1 and Remark 7.2 show that the covering-space result does not imply global isocausality, and one should expect only a one-way causal relation between the compactified models. This distinction between local/covering and global/quotient behavior is essential for understanding the true strength of the causal equivalences in these examples.
@@
In the context of ongoing research on spacetimes with closed timelike curves [1, 4, 9, 19], our findings offer a unified framework to view several causality-violating spacetimes through a common geometric and causal lens. The work reveals the intricate relationships between these examples and motivates a broader investigation into the interplay between geometry, causality, and spacetime topology. For instance, it would be compelling to investigate whether other spacetime models, such as the pseudo-Kerr spacetime [3]—a rotating generalization of the pseudo-Schwarzschild metric—also belong to this family of causality-violating, Misner-type spacetimes.
We view this article as a step toward a broader classification of causality-violating spacetimes. The warped-product ansatz and the explicit isocausality mappings presented here suggest that tools from conformal geometry and causal theory can provide a systematic framework for organising, comparing, and understanding causal structures in a wide range of gravitational models.